QUESTION 1

  1. VC firm X makes a $1 million dollar investment. Ten years later the firm liqudiates its investment for $2 million. The IRR of the investment is over 12%.

    True

    False

1.85185 points   

QUESTION 2

  1. Timothy J. Sturgeon implies that a conscience effort by city planners to replicate the Silicon Valley technology culture in the short run may be wasteful. He argues that Silicon Valley evolved over a long period to become the venture capital high technology innovation hub that it is today.

    True

    False
Answer #1

Note: As per HOMEWORKLIB POLICY, only one question is allowed to answer at a time, so here first question is being answered :

ANSWER: False

EXPLANATION:

future value = present value x (1 + i)^n

$2 = $1 x (1 + n)^10

therefore,

1 + n = $2^(1/10)

1 + n = 1.0717

therefore,

n = 1.0717 - 1

= 0.0717

= 7.17%

THEREFORE,

IRR is less than 12%